You measure 27 textbooks' weights, and find they have a mean weight of 70 grams. Assume the population standard deviation is 5.4 grams. Based on this, construct a 94% confidence interval for the true population mean textbook weight. Assume n is too small to allow use of the normal. Instead use a t statistic.
What is the degrees of freedom?

dof =   

What is the confidence interval. Give your answers as decimals, to two places

< μμ <
Answer #1

df = n - 1 = 27 - 1 = 26

Confidence Interval
X̅ ± t(α/2, n-1) S/√(n)
t(α/2, n-1) = t(0.06 /2, 27- 1 ) = 1.967
70 ± 1.967 * 5.4/√(27)
Lower Limit = 70 - 1.967 * 5.4/√(27)
Lower Limit = 67.96
Upper Limit = 70 + 1.967 * 5.4/√(27)
Upper Limit = 72.04

67.96 < < 72.04